OpenStudy (anonymous):

solve equation by taking square roots: k^2+76 how do i do that?

OpenStudy (anonymous):

This is not an equation. In equations, there must be an equality sign.

OpenStudy (anonymous):

the plus was supposed to be =, sorrry. k^2=76

OpenStudy (anonymous):

if the plus sign was an equals sign: the square root of k^2 is k. so k= sqrt of 76. which is 8.717797887....

OpenStudy (anonymous):

But if you're asking about \(k^2+76=0\) then it has no real solution.

OpenStudy (anonymous):

OK then, NK got the answer for you :D.

OpenStudy (anonymous):

\[k^2=76 \implies k=-\sqrt{76}, k=\sqrt{76}\]

OpenStudy (anonymous):

So two solutions.

OpenStudy (anonymous):

so how would you get the answer for k^2=16? ...

OpenStudy (anonymous):

K can have two values; \(k=4\) or \(k=-4\)

OpenStudy (anonymous):

do you just find the square root of the last number?

OpenStudy (anonymous):

Yep, and take both positive and negative values.

OpenStudy (anonymous):

so x2=21 is 4.58.... or no? :(

OpenStudy (anonymous):

Exactly \(x=4.58\) or \(x=-4.58\) :)